# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

### v3

@@ Abstract
The application of formulas is a fundamental ability of humans when addressing numerical reasoning problems. However, existing numerical reasoning datasets seldom explicitly indicate the formulas employed during the reasoning steps. To bridge this gap, we construct a dataset for formula-based numerical reasoning called FormulaReasoning, which consists of 5,420 reasoning-based questions.  We employ it to conduct evaluations of LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thought methods, and we further explore using retrieval-augmented LLMs provided with an external formula database associated with our dataset. We also experiment with supervised methods where we divide the reasoning process into formula generation, parameter extraction, and numerical calculation, and perform data augmentation. Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaReasoning.
@@ 1 Introduction
Numerical reasoning constitutes one of the significant forms within natural language reasoning (Frieder et al., 2023). The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs (OpenAI, 2023; Touvron et al., 2023; Li et al., 2023c) and specialized datasets (Wang et al., 2017; Dua et al., 2019; Amini et al., 2019; Cobbe et al., 2021a). Current datasets for numerical reasoning typically include simple, commonsense numerical questions that do not reflect the complexity of real-world problems. These datasets have not fully addressed the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process. This issue becomes particularly evident when LLMs meet hallucination (Frieder et al., 2023; Bang et al., 2023). Consequently, one might naturally ask “What knowledge could I use to guide numerical reasoning process?”. Formulas exactly represent such knowledge that has been largely overlooked in research but is frequently utilized in real-life applications.
@@
Take a question from the GSM8K (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?”. This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula. However, in our FormulaReasoning dataset, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object.
@@
To address this gap, we constructed a dataset for numerical reasoning that requires the use of formulas called FormulaReasoning. We annotated formulas for each question in FormulaReasoning. An example of FormulaReasoning is shown in Figure 1.111Please note that FormulaReasoning is in Chinese. For the convenience of understanding, we translated Chinese into English in all the examples presented in this paper. The formula-based feature makes FormulaReasoning a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems. Indeed, in fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge. However, existing datasets scarcely consider explicit incorporation of formulas into numerical reasoning.
Dataset Math23K-F MAWPS-F GSM8K FormulaReasoning # questions 23,162 2,373 8,792 5,420 # formulas (and variants) 51 (131) 18 (46) 0 (0) 272 (824) # questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) N/A 5,420 (100%) Avg. # reasoning steps 1.16 1.01 3.59 2.37
Table 1: Statistics of Math23-F, MAWPS-F, GSM8K and our FormulaReasoning.
We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations. With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and a set of relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database. The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems. In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known GSM8K. In comparison to Math23K-F and MAWPS-F, FormulaReasoning contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas. Additionally, all questions in FormulaReasoning require the use of formulas. The higher average number of reasoning steps (2.37 vs. 1.16/1.01) implies that FormulaReasoning is more challenging and better suited for evaluating existing models as a multi-step formula-based reasoning task.
We used FormulaReasoning to evaluate LLMs ranging from 7B to >100B parameters, as well as fine-tuned models such as Qwen-1.8B (Bai et al., 2023) and ChatGLM-6B (Zeng et al., 2022) with a proposed Chain-of-Thought supervised fine-tuned method and a data augmentation method. We also trained an encoder for formula retrieval and experimented with retrieval-augmented generative models. Our empirical findings show that the best existing models only achieve an accuracy of around 74%, lagging behind an accuracy 92% of humans, indicating that there is still significant room for exploration in formula-based numerical reasoning.
@@
The dataset is available on https://zenodo.org/doi/10.5281/zenodo.11408109 under the CC BY 4.0 License and our code is available on https://github.com/nju-websoft/FormulaReasoning under the Apache License 2.0.
@@ 2.1 Numerical Reasoning Datasets
Numerical reasoning is one of the fundamental capabilities of natural language reasoning. The study of numerical reasoning in natural language has existed for several years. Numerous datasets, such as DROP (Dua et al., 2019), GSM8K (Cobbe et al., 2021b), TSQA (Li et al., 2021) and MATH (Hendrycks et al., 2021), have introduced natural language numerical reasoning. Another line of research focusing on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., a question) and require the generation of an arithmetic expression that can compute an answer. Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc.
@@ 2.2 Numerical Reasoning Methods
We established representative existing methods as baselines for FormulaReasoning, including zero/few-shot CoTs prompting methods to LLMs ranging from 7B to over 100B parameters. We trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning. We also divided the reasoning process into formula generation, parameter extraction, and calculation, and used data augmentation to enhance fine-tuned models with fewer than 7B parameters.
@@ 3 Dataset Construction
We collected raw questions from Chinese junior high school physics examinations from 2015 to the present. We had a total of five postgraduate volunteer students, and they all hold a bachelor’s degree in science and engineering. We then annotated the reasoning steps and corresponding formulas for each question. This process involved a combination of manual annotation and the assistance of LLMs to improve the efficiency of annotation. Each question is associated with an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all the parameters within the formulas. Finally, we compiled all the formulas we merged those expressing the same meaning to create a formula database. We describe this process to construct FormulaReasoning in detail below.
@@ 3.1 Preprocessing
Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical. Utilizing a regular expression-based approach, we extracted the final numeric answer, including its unit, from the explanation. We found that for 487 questions, the regular expressions did not return results, so we manually annotated the positions of their answers in the text explanations. Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
@@ 3.3 Formula Database Construction
Table 3: Changes in the number of formulas after each merging step.
@@
In the symbolic rules based merging process, the semantic information of the parameter names was neglected. This led us to perform merges grounded on the semantics of the parameter names. For instance, two formulas that were not merged during the symbolic fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged. We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, "density" and "density of water" are semantically similar). Specifically, for formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^4] We used jieba: https://github.com/fxsjy/jieba.. When the two sets overlapped, the parameters were considered to have semantic connection, and the formulas became candidates for merging. Utilizing this approach, we identified a set of pairs of potentially mergeable formulas and then consulted the LLM for a thorough evaluation of each pair. The prompts can be found in Appendix A.1.3. After this step, the number of formulas in the formula database was reduced to 439.
@@ 4.2.2 LLMs
We conducted experiments on GPT-4-turbo, GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters. We also evaluated on Llama2-7B (Touvron et al., 2023), Llama3-8B (Meta, 2024), Qwen-7B/14B (Bai et al., 2023), InternLM2-7B/20B (Team, 2023), ChatGLM3-6B (Zeng et al., 2022), including the base and chat versions of these models. We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat or instruct versions.
@@ 4.2.4 Supervised Fine-tuned Models
We found that directly prompting models possessing fewer than 7B parameters failed to produce satisfactory outcomes (for example, ChatGLM3-6B attained merely 8.99 points in a zero-shot setting). Therefore, we conducted supervised fine-tuning of models with fewer than 7B parameters, yet discerned that, dissimilar to larger models (such as GPT-4-turbo), smaller models did not exhibit proficient performance in numerical extraction and calculation. In order to augment the reasoning capabilities of smaller models, we explored two approaches for improvement.
@@
We decomposed the reasoning process into several steps. First, we instructed the model to generate the formulas required to solve the question. Subsequently, the parameter names within the formulas were extracted, allowing the model to retrieve the corresponding values and units from the context. Next, the formulas and the associated parameter values were provided to a calculator to obtain the final result. This approach relieved the model of the numerical calculation, allowing it to concentrate on the reasoning aspect.
@@
We augmented the training dataset with the assistance of larger models. Firstly, we utilized a few-shot approach to prompt the LLM (Qwen-max) to generate new question-answer pairs. The correctness of the computation process generated by the LLM was meticulously verified using a calculator. Subsequently, the formulas generated by the model were extracted and normalized. More details could be found in Appendix A.3.1.
@@ 5 Experiments Results
In this section, we presented the experimental results and analysis. Due to space constraints, the error analysis can be found in Appendix A.2 and the implementation details can be found in Appendix A.4.
@@ 5.2 Results of LLMs
Model Size zero-shot CoT few-shot CoT id test ood test Avg. id test ood test Avg. GPT-4-turbo unknown 70.07 72.89 71.43 71.50 77.49 74.38 GPT-3.5-turbo unknown 26.13 25.58 25.87 32.07 29.92 31.03 GLM4 >100B 65.32 65.22 65.27 62.47 65.98 64.16 Qwen-max >100B 58.67 57.80 58.25 58.91 63.94 61.33 InternLM 20B 5.70 4.60 5.17 18.29 11.25 14.90 Qwen 14B 32.07 37.60 34.73 44.89 36.83 41.01 Llama3 8B 26.66 17.98 20.41 12.81 8.87 10.91 Llama2 7B 0.00 0.26 0.13 1.43 0.26 0.87 Qwen 7B 7.36 8.70 8.01 21.14 18.16 19.71 InternLM 7B 7.84 7.67 7.76 9.50 8.18 8.86 ChatGLM3 6B 9.36 8.62 8.99 23.89 19.95 21.92 Human - 93.49 90.47 92.03 93.49 90.47 92.03
Table 4: Results of LLMs with zero-shot and few-shot prompting. indicates that the chat or instruct version of the model was used in the zero-shot setting, while the base version of the model was used in the few-shot setting.
The evaluation results on LLMs are shown in Table 4. GPT-4-turbo exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked GLM4 by an average of 6.16 points in zero-shot setting and 10.22 in few-shot setting. Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings. The subpar performance of Llama2 might be due to its pre-training data being primarily in English. We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.4.3.  After incorporating few-shot examples, GPT-4-turbo, GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14. However, similar performance changes were not observed on GLM4, possibly due to its supervised fine-tuning and alignment with human preferences which enhanced GLM4’s understanding of instructions but probably also compromised its in-context learning ability.
Human performance surpassed the performance of few-shot GTP-4-turbo on the id and ood tests by margins of 21.99 and 13.25 points, respectively. Such results demonstrated that there remained a substantial gap between the current capabilities of state-of-the-art LLMs and human performance. This was even more pronounced when considering smaller-scale models. These findings underscored the challenging nature of FormulaReasoning as an unresolved dataset, and that there was significant room for improvement in LLMs as they struggled to match human levels of reasoning.

### v4

@@ Abstract
The application of formulas is a fundamental ability of humans when addressing numerical reasoning problems. However, existing numerical reasoning datasets seldom indicate explicitly the formulas employed during the reasoning steps. To bridge this gap, we construct a dataset for formula-based numerical reasoning called FormulaReasoning, which consists of 5,420 reasoning-based questions.  We employ it to conduct evaluations of LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thought methods, and we further explore using retrieval-augmented LLMs provided with an external formula database associated with our dataset. We also experiment with supervised methods where we divide the reasoning process into formula generation, parameter extraction, and numerical calculation, and perform data augmentation. Our empirical findings underscore the significant potential for improvement in existing models when applied to our challenging, formula-driven FormulaReasoning.
@@ 1 Introduction
Numerical reasoning constitutes one of the significant forms within natural language reasoning (Frieder et al., 2023). The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs (OpenAI, 2023; Touvron et al., 2023; Li et al., 2023c) and specialized datasets (Wang et al., 2017; Dua et al., 2019; Amini et al., 2019; Cobbe et al., 2021a). Current datasets for numerical reasoning typically include simple, commonsense numerical questions that do not reflect the complexity of real-world problems. These datasets have not fully addressed the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process. This issue becomes particularly evident when LLMs meet hallucination (Frieder et al., 2023; Bang et al., 2023). Consequently, one might naturally ask “What knowledge could be used to guide numerical reasoning process?”. Formulas exactly represent such knowledge that has been largely overlooked in previous research but is frequently utilized in real-life applications.
@@
Take a question from the GSM8K (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?”. This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula. However, in our FormulaReasoning dataset, we require domain-specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object.
@@
To fill this gap, we constructed a dataset for numerical reasoning that requires the use of formulas called FormulaReasoning. We annotated formulas for each question in FormulaReasoning. An example of FormulaReasoning is shown in Figure 1.111Please note that FormulaReasoning is originally in Chinese. For the convenience of understanding, we translated Chinese into English in all the examples presented in this paper. The formula-based feature makes FormulaReasoning a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems. Indeed, in fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge. However, existing datasets scarcely consider explicit incorporation of formulas into numerical reasoning.
Table 1: Statistics of Math23-F, MAWPS-F, GSM8K, MATH and our FormulaReasoning.
Dataset Math23K-F MAWPS-F GSM8K MATH FormulaReasoning # questions 23,162 2,373 8,792 12,500 5,420 # formulas (and variants) 51 (131) 18 (46) 0 (0) 0 (0) 272 (824) # questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) N/A N/A 5,420 (100%) Avg. # reasoning steps 1.16 1.01 3.59 Not Provided 2.37
We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations. With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and a set of relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a consolidated formula database. The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems. In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known GSM8K and MATH (Hendrycks et al., 2021). In comparison to Math23K-F and MAWPS-F, FormulaReasoning contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas. Additionally, all questions in FormulaReasoning require the use of formulas. The higher average number of reasoning steps (2.37 vs. 1.16/1.01) implies that FormulaReasoning is more challenging and better suited for evaluating existing models as a multi-step formula-based reasoning task.
We used FormulaReasoning to evaluate LLMs ranging from 7B to $>$100B parameters, as well as fine-tuned models such as Qwen-1.8B (Bai et al., 2023) and ChatGLM3-6B (Zeng et al., 2022) with a proposed Chain-of-Thought supervised fine-tuned method and a data augmentation method. We also trained an encoder for formula retrieval and experimented with retrieval-augmented generative models. Our empirical findings show that the best existing models only achieve an accuracy of around 84%, lagging behind an accuracy 92% of humans, indicating that there is still significant room for exploration in formula-based numerical reasoning.
@@
The dataset is available on Zenodo [^2] https://zenodo.org/doi/10.5281/zenodo.11408109 under the CC BY 4.0 License and our code is available on GitHub [^3] https://github.com/nju-websoft/FormulaReasoning under the Apache License 2.0.
@@ 2.1 Numerical Reasoning Datasets
Numerical reasoning is one of the fundamental capabilities of natural language reasoning. The study of numerical reasoning in natural language has existed for several years. Numerous datasets, such as DROP (Dua et al., 2019), GSM8K (Cobbe et al., 2021b), TSQA (Li et al., 2021) and MATH (Hendrycks et al., 2021), have introduced natural language numerical reasoning. Another line of research focusing on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., a question) and require the generation of an arithmetic expression that can compute an answer. Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc. Several works focus on specialized domains where some of the questions in their datasets require numerical reasoning. Examples include GeoSQA (Huang et al., 2019), which focuses on the geography domain, the STEM (Drori et al., 2023) dataset and the ScienceQA (Lu et al., 2022) which covers multiple disciplines in science and technology. The distinguishing feature of our FormulaReasoning is that the numerical reasoning questions within these datasets lack explicitly labeled formulas.
@@ 2.2 Numerical Reasoning Methods
We established representative existing methods as baselines for FormulaReasoning, including zero/few-shot CoTs prompting methods to LLMs ranging from 7B to over 100B parameters. We trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning. We also divided the reasoning process into formula generation, parameter extraction, and numerical calculation, and used data augmentation to enhance fine-tuned models with fewer than 7B parameters.
@@ 3 Dataset Construction
We collected raw questions from Chinese junior high school physics examinations from 2015 to the present. We had a total of five postgraduate volunteer students, and they all hold a bachelor’s degree in science and engineering. We then annotated the reasoning steps and corresponding formulas for each question. This process involved a combination of manual annotation and the assistance of LLMs to improve the efficiency of annotation. Each question is associated with an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all the parameters within the formulas. Finally, we compiled all the formulas and we merged those expressing the same meaning to create a formula database. We describe this process to construct FormulaReasoning in detail below.
@@ 3.1 Preprocessing
Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical. Utilizing a regular expression-based approach, we extracted the final numerical answer, including its unit, from the explanation. We found that for 487 questions, the regular expressions did not return results, so we manually annotated the positions of their answers in the text explanations. Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
@@ 3.3 Formula Database Construction
Table 3: Changes in the number of formulas after each merging step.
@@
In the symbolic rules based merging process, the semantic information of the parameter names was neglected. This led us to perform merges grounded on the semantics of the parameter names. For instance, two formulas that were not merged during the symbolic fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged. We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, ”density” and ”density of water” are semantically similar). Specifically, for formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^6] We used jieba: https://github.com/fxsjy/jieba.. When the two sets overlapped, the parameters were considered to have semantic connection, and the formulas became candidates for merging. Utilizing this approach, we identified a set of pairs of potentially mergeable formulas and then consulted the LLM for a thorough evaluation of each pair. The prompts can be found in Appendix A.1.3. After this step, the number of formulas in the formula database was reduced to 439.
@@ 4.2.2 LLMs
We conducted experiments on GPT-4o, GPT-4-turbo, GPT-3.5-turbo, GLM-4-plus, GLM-4-flash (GLM et al., 2024), and Qwen-max. We also evaluated on Qwen2.5-7B/14B (Yang et al., 2024) and Llama3.1-8B (Meta, 2024).
@@ 4.2.4 Supervised Fine-tuned Models
We found that directly prompting models possessing fewer than 7B parameters failed to produce satisfactory outcomes (for example, ChatGLM3-6B attained merely 8.99 points in a zero-shot setting). Therefore, we conducted supervised fine-tuning of models with fewer than 7B parameters, yet discerned that, dissimilar to larger models (such as GLM-4-plus), smaller models did not exhibit proficient performance in numerical extraction and calculation. In order to augment the reasoning capabilities of smaller models, we explored two approaches for improvement. We conducted experiments on Qwen-1.8B (Bai et al., 2023) and ChatGLM3-6B (Zeng et al., 2022).
@@
We decomposed the reasoning process into several steps. First, we instructed the model to generate the formulas required to solve the question. Subsequently, the parameter names within the formulas were extracted, allowing the model to retrieve the corresponding values and units from the context. Next, the formulas and the associated parameter values were provided to a calculator to obtain the final result. This approach relieved the model from numerical calculation, allowing it to concentrate on the reasoning aspect.
@@
We augmented the training dataset with the assistance of larger models. Firstly, we utilized a few-shot approach to prompt a LLM (Qwen-max) to generate new question-answer pairs. The correctness of the computation process generated by the LLM was meticulously verified using a calculator. Subsequently, the formulas generated by the model were extracted and normalized. More details could be found in Appendix A.2.1.
@@ 4.3 Metric
4.4 Implementation Details
We accessed to GPT-4o (gpt-4o-2024-08-06 version), GPT-4-turbo (gpt-4-1106-preview version), GPT-3.5-turbo (gpt-3.5-turbo-1106 version) [^7] https://platform.openai.com/docs, GLM-4-plus, GLM-4-flash [^8] https://open.bigmodel.cn/, Qwen-max and Qwen2.5-7B/14B [^9] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters. For Llama3.1, we conducted experiments on NVIDIA V100-32G GPUs. These LLMs generated using nucleus sampling with top_p=0.8. Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0. For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs. For ChatGLM3-6B, we fine-tuned with LoRA (Hu et al., 2021) with r=8, alpha=32 and learning rate of 5e-5, batch size of 1. The max input length and output length are both set to 512. We utilized nucleus sampling with top_p=0.8 for generation. In the case of CoT-SFT, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated. Training Qwen-1.8B, ChatGLM3-6B models required 12 and 24 hours respectively.
@@ 5.2 Results of LLMs
Table 4: Results of LLMs with zero-shot and few-shot prompting.
Model Size zero-shot CoT few-shot CoT id test ood test Avg. id test ood test Avg. GPT-4o unknown 77.20 72.38 74.88 76.01 73.66 74.88 GPT-4-turbo unknown 70.07 72.89 71.43 71.50 77.49 74.38 GPT-3.5-turbo unknown 26.13 25.58 25.87 32.07 29.92 31.03 GLM-4-plus $>$100B 84.32 81.07 82.76 82.90 85.68 84.24 GLM-4-flash unknown 71.50 71.87 71.68 61.76 67.01 64.29 Qwen-max $>$100B 57.24 60.10 58.62 55.82 61.38 58.50 Qwen2.5 14B 61.28 64.71 62.93 61.28 65.22 63.18 Qwen2.5 7B 42.04 43.73 42.38 59.62 65.73 62.56 Llama3.1 8B 13.06 9.74 11.46 9.74 9.72 9.73 Human - 93.49 90.47 92.03 93.49 90.47 92.03
The evaluation results on LLMs are shown in Table 4. GLM-4-plus exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked GPT-4o by an average of 7.88 points in zero-shot setting and 9.36 in few-shot setting. Among models with size not exceeding 20B, Qwen2.5-14B demonstrated commendable performance in both zero-shot and few-shot settings. The subpar performance of Llama3.1 might be due to its pre-training data being primarily in English.  After incorporating few-shot examples, GPT-4-turbo, GPT-3.5-turbo, GLM-4-plus and Qwen2.5 demonstrated performance improvements, ranging from 0.25 to 20.18. However, similar performance changes were not observed on other LLMs. Surprisingly, the open-source Qwen2.5-14B model outperformed the closed-source Qwen-max model [^10] We have not yet found clear information indicating whether the closed-source Qwen-max is also based on version 2.5..
Human performance surpassed the performance of the flagship model GLM-4-plus with zero-shot setting and few-shot setting by margins of 9.27 and 7.79 points, respectively. Such results demonstrated that there remained a substantial gap between the current capabilities of state-of-the-art LLMs and human performance. This was even more pronounced when considering smaller-scale models. These findings underscored the challenging nature of FormulaReasoning as an unresolved dataset, and that there was significant room for improvement in LLMs as they struggled to match human levels of reasoning.
We also compared the chain of thought (CoT) and program of thought (PoT, Chen et al., 2023) methods, with the results presented in Appendix A.2.4. The results indicated that CoT and PoT demonstrated varying performances between different models and under different settings.